OpenStudy (anonymous):

Cashews cost $5 per pound and peanuts cost $1.50 per pound. How many pounds of cashews should be mixed with peanuts to get a mixture of 32 pounds that will cost $4 per pound? Round to the nearest tenth.

OpenStudy (mertsj):

c = # pounds of cashews p = # pounds of peanuts p+c=32 5c+1.5p=4(32)